# Self-Filtered Distillation with LLMs-generated Trust Indicators for Reliable Patent Classification

### v1

@@ 5 Conclusion
This paper proposes SFD, a framework that improves reliability and accuracy in LLM-based patent classification by leveraging reasoning quality as a supervision signal. Unlike prior methods relying solely on model outputs, our approach evaluates rationales with trust metrics and incorporates them into training. Three trust metrics comprise Stability Consistency, Class Entailment Alignment, and Logic Agreement Score. Their aggregated scores dynamically weight or filter training samples, allowing the model to emphasize high-quality rationales while suppressing noisy ones.
SFD leads to threefold contributions: (1) establishing a protocol for assessing rationale trustworthiness, (2) designing a trust-driven weighting and filtering strategy, and (3) demonstrating consistent gains across LLM backbones and patent benchmarks. Experimental results show that trust-aware supervision not only enhances predictive performance and interpretability, but also establishes a new paradigm for reliable LLM-based reasoning in high-stakes patent analytics. As for future work, we will extend the SFD framework to other high-stakes domains such as biomedical and legal texts.
@@
This study is conducted on the publicly available USPTO-2M dataset, and thus does not involve issues related to copyright or personal data. However, the proposed trust-based learning framework is designed for high-stakes domains such as patents, where misinterpretation of model-generated or model-evaluated reasoning could potentially influence legal or industrial decision-making. Therefore, the methods presented in this work are intended for research purposes, and any practical deployment should be accompanied by careful expert validation.
@@
For transparency, we note that ChatGPT was used to improve grammar and clarity of expression.

### v3

@@ 5 Conclusion
This study introduces Self-Filtered Distillation (SFD), a framework utilizing three unsupervised trust metrics to mitigate noise in LLM-generated rationales. Empirical evaluations on the USPTO-2M dataset show that SFD significantly outperforms conventional distillation, achieving up to a 38.7% relative improvement in Macro-F1. We confirmed the framework’s statistical robustness through five independent runs and its model-agnostic nature across diverse student architectures. The strong correlation with human judgments ($\rho=0.685$) validates that our trust indicators effectively capture reasoning quality. Ultimately, SFD establishes a reliable paradigm for transferring LLM reasoning to specialized domains like patent classification, ensuring both high performance and stability.
@@
This study is conducted on the publicly available USPTO-2M dataset and thus does not involve issues related to copyright or personal data. However, the proposed trust-based learning framework is designed for high-stakes domains such as patents, where misinterpretation of model-generated or model-evaluated reasoning could potentially influence legal or industrial decision-making. Therefore, the methods presented in this work are intended for research purposes, and any practical deployment should be accompanied by careful expert validation.
@@
For transparency, we note that ChatGPT and Gemini was used to improve grammar, clarity of expression and help to generated figures.